OpenStudy (anonymous):

To break even in a manufacturing business, income or revenue R must equal the cost of production C. The cost C to produce x skateboards is C = 156 + 18x. The skateboards are sold wholesale for $24 each, so revenue R is given by R = 24x. Find how many skateboards the manufacturer needs to produce and sell to break even. (Hint: Set the cost of expression equal to the revenue expression and solve for x.)

OpenStudy (anonymous):

The hint has basically told you what do to to solve the problem. :) C= 156 +18x R= 24x Then to break even, what condition must be met? Cost must equal revenue, so C = R. Substitute and solve.

OpenStudy (anonymous):

x = 6?

OpenStudy (anonymous):

No, not quite. Here is what you have: C = R 156 + 18x = 24x Can you solve for x?

OpenStudy (anonymous):

Oops, x is 26 ..

OpenStudy (anonymous):

Si. :D

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

^^ Welcome.